# Metric theory of inhomogeneous Diophantine approximations with a fixed matrix

### v2

@@ Abstract
In this paper we develop a metric theory of inhomogeneous Diophantine approximation for the case of a fixed matrix. We use transference principle to connect uniform Diophantine properties of a pair $(\Theta,\boldsymbol{\eta})$ of a matrix and a vector with the asymptotic Diophantine properties of the transposed matrix $\Theta^{\top}$, and vice versa, the asymptotic Diophantine properties of a pair $(\Theta,\boldsymbol{\eta})$ with asymptotic Diophantine properties of the transposed matrix. In these setups, we prove analogues of classical statements of metrical homogeneous Diophantine approximations and answer some open questions that were raised in recent works.
@@ 1.1 Setup and notation
Throughout this paper $m$ and $n$ are positive integers and $d=m+n$. Let ${\bf M}_{n,m}$ be the set of all real $n\times m$ matrices. Let
@@
be a real $n\times m$ matrix and a real vector. We use the notation $|\cdot|$ for the supremum norm of a vector, and the notation $||\cdot||_{\mathbb{Z}^{n}}$ will be reserved for the distance to the nearest integer vector function; that is, for ${\bf v}\in\mathbb{R}^{n}$ we set
@@
In both uniform and asymptotic cases, we want to compare the form $\Theta{\bf q}-\boldsymbol{\eta}||_{\mathbb{Z}^{n}}$ with the values of certain functions. Thus, it is essential to introduce the notions of Dirichlet and approximable pairs relative to a certain function. We are following the standard notation from recent works, see [36, 37].
@@
$f_{1}(T):=\frac{1}{T}\,\,\,\,\,\,\,\,\,\,\,\,\text{and, in general,}\,\,\,\,\,\,\,\,\,\,\,\,f_{k}(T):=\frac{1}{T^{k}}.$ (1)
@@
From now on, we will assume that $f$ is a decreasing function with an inverse function $f^{-1}$. To formulate our results we should consider function $g$ defined by
$g(T)=\frac{1}{f^{-1}\left(\frac{1}{T}\right)}.$ (2)
@@
$f(T)=\frac{1}{g^{-1}\left(\frac{1}{T}\right)}.$ (3)
For a function $f$ and constant $C$, we will use the notation $f\circ C$ to denote the composition of multiplication by $C$ with $f$, that is, $\left(f\circ C\right)(T)=f(CT)$.
@@ 1.2 Selected results
We would like to start with formulating two theorems, which provide special cases of some selected results of this paper. We formulate general versions of these statements, as well as some other related statements, in Section 3. Note that part A of Theorem 3 and part A of Theorem 4 follow from arguments due to Jarník, which we include for the complete picture; see their slightly different forms in Section 3.
We begin with the uniform case. The sets $\widehat{\bf D}[g]$, as well as their $\boldsymbol{\eta}$-fixed sections $\{\Theta:\,\,\,(\Theta,\boldsymbol{\eta})\in\widehat{\bf D}[g]\}$ were actively studied before (see [37] for a metric criterion for pairs, [32] for generalizations on Hausdorff measures, [1] for dimension bounds for sections). The recent paper [1] by Aggarwal also provides the upper bounds (Theorem 1.8) for the Hausdorff dimension of the sets $\widehat{\bf D}^{\Theta}=\widehat{\bf D}^{\Theta}[f_{1}]$ (more precisely, they work with bigger sets of divergent in average trajectories). Less was known about the sets $\widehat{\bf D}^{\Theta}[g]$ for general $g$. Our first theorem describes some metric properties of these sets; in particular, we give answers to the questions stated in Section 7.3 in [37] and Section 5.1 in [32]. For more general version of Theorem 3 B, see Theorem 1. Define
$\kappa=2^{1-m-n}((m+n)!)^{2}.$
@@
We note that Theorem 3 implies the recent result from [37] (showing that the latter is equivalent to the Khintchine-Groshev’s theorem); see Section 4.4 for some discussion and deduction.
The second theorem answers some metrical questions about the sets $\widehat{\bf W}_{n,m}^{\Theta}\left[g\right]$. The statement of Theorem 4, part B can be thought of as a direct analog to the fact that "the set of badly approximable objects has zero measure" (cf. [5]), where part A serves as a reference universal bound to compare with; see Theorem 3 for the language, corresponding to this description, and Theorem 4 for the winning properties of the corresponding sets.
Part C utilizes the notation $Y_{\nu}=|{\bf y}_{\nu}$| for the lengths of the best approximation vectors for $\Theta^{\top}$; we will introduce all the necessary definitions in Section 5.2. This statement provides a partial convergence condition for functions $g$ satisfying several technical restrictions; for the general version of this statement, see Theorem 6. We note that the optimal convergence-divergence condition for these sets is only known in case $n=m=1$ (see [18]), and the proof relies on the fact that there are "no nontrivial singular numbers".
@@
  3. C.
    Let $\lambda$ be a decreasing function on $\mathbb{R}_{+}$ such that (8) holds. Let $g$ be such that the conditions (7), (11) and (12) hold.
@@
    $\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (4)
@@
The asymptotic (ordinary) setting also begins with the argument by Jarník (Theorem B), which provides an analog of Dirichlet’s theorem corollary. We show that that the generalization of bad approximability provides us with the set of measure zero, thus completing the result from [31]. We also show that this set of badly approximable $\boldsymbol{\eta}$ is winning in Hyperplane absolute game. Finally, we show a sufficient condition for a Khintchine-Groshev type theorem, showing that a set of shift vectors $\boldsymbol{\eta}$ providing us with well enough approximable pairs $(\Theta,\eta)$ has measure zero.
In addition, we revisit some classical results by Khintchine regarding the connection between ordinary homogeneous and uniform inhomogeneous approximation. In particular, Khintchine’s approach gives an easy criterion of badly approximablity (Corollary 9).
Our central technical tool is the classical transference principle. An extensive theory of transference, that is connection between inhomogeneous approximations to matrix $\Theta$ and homogeneous approximations ($\boldsymbol{\eta}=\boldsymbol{0}$) to transposed matrix $\Theta^{\top}$ was developed in classical works by A. Khintchine [26, 27, 28, 29] and V. Jarník [23, 24, 25]. A detailed exposition of the basic moments of this theory one can find in Chapter V of the book [11] by J.W.S. Cassels.
@@ 1.3 Structure of the paper
The structure of our paper is as follows. In Section 2 we introduce all necessary notions and notation. Here we should note that many of notation from recent papers differ from those from Cassels’ book [11] and earlier works. To make our exposition more clear we tried everywhere in the text to explain the meaning of modern results and their generalizations in terms of classical notation and compare the corresponding results.
@@
Section 3.1 deals with uniform inhomogeneous approximation in connection with asymptotic approximation for transposed matrix $\Theta^{\top}$.
Section 3.2 is devoted to asymptotic (or ordinary) inhomogeneous approximation and its connections with uniform approximation for transposed matrix $\Theta^{\top}$.
@@
Section 4 shows some applications and corollaries of our results; specifically, in Section 4.1 we show what our main theorems say applied to functions of form $T^{a}\left(\log T\right)^{b}$. In Section 4.2 some old and recent characterizations of sets of badly and very well approximable matrices are discussed. In Section 4.3 we recall and slightly generalize the inequalities on Diophantine exponents due to Bugeaud and Laurent. In Section 4.4 we show how our results provide a simple proof of the Dirichlet theorem for pairs by D. Kleinbock and N. Wadleigh.
Proofs of the results from Section 3.1 and Section 3.2, together with reformulations on the language of irrationality measure functions, will be given in Section 7. Specifically, in Section 7.1 for Theorem 2, Theorem 4, Theorem 6, and in addition Theorem A and Theorem B; in Section 7.2 for Theorem 1 and in Section 7.3 for Theorem 3).
@@
The remaining sections are technical. In Section 5 we introduce the language of best approximation vectors and irrationality measure functions which will be used for proofs, and which allows one to compare modern and classical results. Specifically, in Section 5.2 we prove a simple property of irrationality measure function that may be of a separate interest, and Section 5.3 provides a short discussion about asymptotic directions of the best approximations vectors, which will be important for some theorems in section 3.
Section 6 collects some technical statements which allow us to generalize our results to certain affine subspaces: we introduce the sets of exceptions for our theorem in Section 6.1 and prove technical metrical lemmata in Section 6.2.
Section 8 shows how Theorem 7 and Theorem 4 C are deduced from Theorem 6.
@@ 2.1.1 Bad and very good approximability
The definition of $g$-approximable matrices given above motivates a natural generalization of the notion of Badly approximable pairs. The classical definition (see [33] for precise formulations) says that a pair is badly approximable if, up to a constant, this pair can not be asymptotically approximated by a function decreasing faster than $f_{1}$. One can replace $f_{1}$ by an arbitrary function $g$ and obtain
@@
We will prove that under certain choices of $g$ the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is a Lebesgue null set (Section 3.2.2), but its restrictions on certain subspaces are winning in variations of Schmidt’s game (Section 3.2.3). We will also discuss some necessary and sufficient conditions for a matrix to be badly approximable in classical sense (that is, belong to the set ${{\bf BA}}_{n,m}\left[f_{1}\right]$) in Section 4.2 (where $f_{1}$ is defined in (1)).
Another notion we will briefly discuss in Section 4.2 is very well approximability, which we will only formulate for matrices:
@@ 2.1.2 Singularity
Analogously to the classical homogeneous case, one can call the elements of $\widehat{\bf D}_{n,m}^{\Theta}[\varepsilon g]$ for $0<\varepsilon<1$ $g$-Dirichlet improvable shift vectors. This suggests a natural definition of singularity:
@@
We will call a pair $(\Theta,\boldsymbol{\eta})$ simply singular if it is $f_{1}$-singular with $f_{1}$ defined in (1) and denote the set of all singular pairs $\widehat{\bf Sing}_{n,m}$.
@@
$\left(\begin{array}[]{c}\theta_{1,1}\cr\vdots\cr\theta_{n,1}\end{array}\right),\dots,\left(\begin{array}[]{c}\theta_{1,m}\cr\vdots\cr\theta_{n,m}\end{array}\right),\left(\begin{array}[]{c}1\cr\vdots\cr 0\end{array}\right),\dots,\left(\begin{array}[]{c}0\cr\vdots\cr 1\end{array}\right)$ (5)
@@ 2.1.3 Diophantine exponents
has solutions ${\bf q}\in\mathbb{Z}^{m}\setminus\{{\bf 0}\}$ for any $t\in\mathbb{R}_{+}$ large enough. We define the corresponding ordinary and uniform Diophantine exponents for matrix $\Theta$ by
@@ 2.1.4 HAW property
Lastly, we briefly discuss the Hyperplane absolute winning game and HAW property of the sets.
Hyperplane absolute winning game is a version of the classical two-player Schmidt’s game (see [45] for the Schmidt’s $(\alpha,\beta)$-game setup) introduced by Broderick, Fishman, Kleinbock, Reich and Weiss in [9], which provides the notion of winning stronger than the original Schmidt’s one. We call the set Hyperplane absolute winning (HAW) if it is winning in this game. We will not need the definitions and rules of the game and thus will not define it here, just mentioning several simple properties of HAW sets which are of interest for this paper:
@@
  • Countable intersection of HAW sets is HAW.
@@ 2.2 Technical conditions on functions $f$ and $g$
Property (6) is natural: it is equivalent to the existence of such $K>1$ that
@@
For instance, it is satisfied by all the functions we use in Section 4. For $f$ and $g$ connected by (2), this is equivalent to saying that
$g(CT)\geq C^{-\frac{1}{\gamma}}g(T)\,\,\,\,\text{for any}\,\,T>1\,\,\text{and for any}\,\,C\,\,\text{large enough}.$ (7)
The remaining stricter conditions are only used in Theorem 7 and Theorem 4 C.
@@
$\text{There exists}\,\,\alpha>0,\,\,\text{such that}\,\,\lambda(T^{\delta})\geq\delta^{-\alpha}\lambda(T)\,\,\text{for any}\,\,\delta>1,\,\,T>0.$ (8)
@@
$\frac{f(T_{2})}{f(T_{1})}<\left(\frac{\log T_{1}}{\log T_{2}}\right)^{\alpha}\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,T_{2}>T_{1},$ (9)
where $\alpha$ is as in (8) (which says that $f$ locally decreases not too slowly), and
$f(T)\geq T^{-\beta}\,\,\,\,\,\,\,\,\ \text{for some}\,\,\beta>0$ (10)
@@
Finally, below are the similar conditions needed in Theorem 4 C.
$\frac{g(T_{2})}{g(T_{1})}<\left(\frac{\log T_{1}}{\log T_{2}}\right)^{\frac{n}{m}\alpha}\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,T_{2}>T_{1},$ (11)
where $\alpha$ is as in (8) and
$g(T)\leq T^{-\frac{1}{\beta}}\,\,\,\,\,\,\,\,\ \text{for some}\,\,\beta>0.$ (12)

### v3

@@ Abstract
In this paper we develop the metric theory of inhomogeneous Diophantine approximation for the case of a fixed matrix. We use transference principle to connect uniform Diophantine properties of a pair $(\Theta,\boldsymbol{\eta})$ of a matrix and a vector with the asymptotic Diophantine properties of the transposed matrix $\Theta^{\top}$, and vice versa, the asymptotic Diophantine properties of a pair $(\Theta,\boldsymbol{\eta})$ with uniform Diophantine properties of the transposed matrix. In these setups, we prove analogues of classical statements of metric homogeneous Diophantine approximations and answer some open questions that were raised in recent works.
Mathematics Subject Classification (2020): 11J20 (Primary), 11J83, 11J13 (Secondary).
@@ 1.1 Setup and notation
Throughout this paper $m$ and $n$ are positive integers. Let ${\bf M}_{n,m}$ be the set of all real $n\times m$ matrices. Let
@@
be a real $n\times m$ matrix and a real vector. We use the notation $|\cdot|$ for the supremum norm of a vector, and the notation $||\cdot||_{\mathbb{Z}^{n}}$ will be reserved for the distance to the nearest integer vector; that is, for ${\bf v}\in\mathbb{R}^{n}$ we set
@@
We want to compare values of the form $\Theta{\bf q}-\boldsymbol{\eta}||_{\mathbb{Z}^{n}}$ with the values of certain functions in both uniform ("system of inequalities is solvable for all parameters large enough") and asymptotic ("system of inequalities is solvable for an unbounded set of parameters") setups. Thus, it is essential to introduce the notions of Dirichlet and approximable pairs relative to a certain function. We are following the standard notation from recent works, see [38, 39].
@@
$f_{1}(T):=\frac{1}{T}\,\,\,\,\,\,\,\,\,\,\,\,\text{and, in general,}\,\,\,\,\,\,\,\,\,\,\,\,f_{k}(T):=\frac{1}{T^{k}}.$ (1.1)
@@
From now on, we will assume that $f$ is a continuous decreasing function with an inverse function $f^{-1}$. To formulate our results we consider the function $g$ defined by
$g(T)=\frac{1}{f^{-1}\left(\frac{1}{T}\right)}.$ (1.2)
@@
$f(T)=\frac{1}{g^{-1}\left(\frac{1}{T}\right)}.$
For a function $f$ and a constant $C$, we will use the notation $f\circ C$ to denote the composition of multiplication by $C$ with $f$, that is, $\left(f\circ C\right)(T)=f(CT)$.
@@ 1.2 Selected results
We would like to start with formulating two theorems, which provide special cases of a few selected results of this paper. We formulate general versions of these statements, as well as some other related statements, in Section 3. Note that part (a) of Theorem 3 and part (a) of Theorem 4 follow from arguments due to Jarník, which we include for the sake of completeness; see their reformulations in Section 3.
We begin with the uniform case. The sets $\widehat{\bf D}[g]$, as well as their $\boldsymbol{\eta}$-fixed sections $\{\Theta:\,\,\,(\Theta,\boldsymbol{\eta})\in\widehat{\bf D}[g]\}$, were actively studied before (see [39] for a metric criterion for pairs, [34] for generalizations on Hausdorff measures, [1] for dimension bounds for sections). The recent paper [1] by Aggarwal also provides the upper bounds (Theorem 1.8) for the Hausdorff dimension of the sets $\widehat{\bf D}^{\Theta}=\widehat{\bf D}^{\Theta}[f_{1}]$ (more precisely, they work with bigger sets of trajectories divergent in average). Less was known about the sets $\widehat{\bf D}^{\Theta}[g]$ for general $g$. Our first theorem describes some metric properties of these sets; in particular, we give answers to the questions stated in Section 7.3 in [39] and Section 5.1 in [34]. For more general version of Theorem 3 (b), see Theorem 1.
Define
$\kappa=2^{1-m-n}((m+n)!)^{2}.$ (1.3)
@@
We note that Theorem 3 implies a recent result from [39] (showing that the latter is equivalent to the Khintchine-Groshev’s theorem); see Section 4.4 for some discussion and a deduction.
The second theorem answers some metrical questions about the sets $\widehat{\bf W}_{n,m}^{\Theta}\left[g\right]$. The statement of Theorem 4, part (b) can be thought of as a direct analog to the fact that "the set of badly approximable objects has zero measure" (cf. [5]), where part (a) serves as a reference point to compare with; see Theorem 3 for the language, corresponding to this description, and Theorem 4 for the winning properties of the corresponding sets.
Part (c) utilizes the notation $Y_{\nu}=|{\bf y}_{\nu}$| for the lengths of the best approximation vectors for $\Theta^{\top}$; we will introduce all the necessary definitions in Section 5.2. This statement provides a partial convergence condition for functions $g$ satisfying several technical restrictions; for the general version of this statement, see Theorem 6. We note that the optimal convergence-divergence condition for these sets is only known in case $n=m=1$ (see [20]), and the proof relies on the fact that there are "no nontrivial singular numbers".
@@
  3. (c)
    Let $\lambda$ be a decreasing function on $\mathbb{R}_{+}$ such that (2.4) holds. Let $g$ be such that conditions (2.3), (2.7) and (2.8) hold.
@@
    $\sum\limits_{\nu=1}^{\infty}\lambda^{\frac{1}{m}}\left(Y_{\nu}^{n}\right)<\infty,$ (1.4)
@@
The asymptotic (ordinary) setting also begins with the argument by Jarník (Theorem B), which provides an analog of Dirichlet’s theorem asymptotic version. We show that that the generalization of bad approximability provides us with the set of measure zero, thus completing the result from [33]. We also show that this set of badly approximable $\boldsymbol{\eta}$ is winning in the Hyperplane absolute game. Finally, we show a sufficient condition for a Khintchine-Groshev type theorem, showing that a set of shift vectors $\boldsymbol{\eta}$ providing us with well enough approximable pairs $(\Theta,\eta)$ has measure zero.
In addition, we revisit some classical results by Khintchine regarding the connection between ordinary homogeneous and uniform inhomogeneous approximation. In particular, Khintchine’s approach gives an easy criterion for badly approximablity (Corollary 9).
Our central technical tool is the classical transference principle. An extensive theory of transference, that is, a connection between inhomogeneous approximations to matrix $\Theta$ and homogeneous approximations ($\boldsymbol{\eta}=\boldsymbol{0}$) to the transposed matrix $\Theta^{\top}$, was developed in classical works by A. Khintchine [28, 29, 30, 31] and V. Jarník [25, 26, 27]. A detailed exposition of the basics of this theory can be found in Chapter V of the book [11] by J.W.S. Cassels.
@@ 1.3 Structure of the paper
The structure of our paper is as follows. In Section 2 we introduce all the necessary notions and notation. Here we should note that some notation in recent papers differs from that in Cassels’ book [11] and earlier works. To make our exposition clearer we tried everywhere in the text to explain the meaning of modern results and their generalizations in terms of classical notation and compare the corresponding results.
@@
Section 3.1 deals with uniform inhomogeneous approximation in connection with asymptotic approximation for the transposed matrix $\Theta^{\top}$.
Section 3.2 is devoted to asymptotic (or ordinary) inhomogeneous approximation and its connections with uniform approximation for $\Theta^{\top}$.
@@
Section 4 shows some applications and corollaries of our results; specifically, in Section 4.1 we show what our main theorems say when applied to functions of form $T^{a}\left(\log T\right)^{b}$. In Section 4.2 some old and recent characterizations of sets of badly and very well approximable matrices are discussed. In Section 4.3 we recall and slightly generalize the inequalities on Diophantine exponents due to Bugeaud and Laurent. In Section 4.4 we show how our results provide a simple proof of the zero-one law for pairs by D. Kleinbock and N. Wadleigh.
Proofs of results from Section 3.1 and Section 3.2, together with reformulations in the language of irrationality measure functions, will be given in Section 7. Specifically, we prove Theorem 2, Theorem 4, Theorem 6, and in addition Theorem A and Theorem B in Section 7.1; we prove Theorem 1 in Section 7.2 and Theorem 3 in Section 7.3.
@@
The remaining sections are technical. In Section 5 we introduce the language of best approximation vectors and irrationality measure functions which will be used for proofs, and which allows one to compare modern and classical results. Specifically, in Section 5.2 we prove a simple property of the irrationality measure function that may be of separate interest, and Section 5.3 provides a short discussion about asymptotic directions of the best approximations vectors, which will be important for some theorems in Section 3.
Section 6 collects some technical statements which allow us to generalize our results to certain affine subspaces: we introduce the sets of exceptions to our theorems in Section 6.1 and prove some technical metrical lemmata in Section 6.2.
Section 8 shows how Theorem 7 and Theorem 4 (c) are deduced from Theorem 6.
Acknowledgements
@@ 2.1.1 Bad and very good approximability
The definition of $g$-approximable matrices given above motivates a natural generalization of the notion of badly approximable pairs. The classical definition (see [35] for a precise formulation) says that a pair is badly approximable if, up to a constant, this pair can not be asymptotically approximated by a function decreasing faster than $f_{1}$. One can replace $f_{1}$ by an arbitrary function $g$ and obtain
@@
We will prove that under certain choices of $g$ the set $\widehat{{\bf BA}}_{n,m}^{\Theta}\left[g\right]$ is a Lebesgue null set (Section 3.2.2), but its restrictions on certain subspaces are winning in variations of Schmidt’s game (Section 3.2.3). We will also discuss some necessary and sufficient conditions for a matrix to be badly approximable in the classical sense (that is, belong to the set ${{\bf BA}}_{n,m}\left[f_{1}\right]$) in Section 4.2 (where $f_{1}$ is defined in (1.1)).
Another notion we will briefly discuss in Section 4.2 is very good approximability, which we will only formulate for matrices:
@@ 2.1.2 Singularity
Analogously to the classical homogeneous case, one can naturally define the notion of singularity:
@@
We will call a pair $(\Theta,\boldsymbol{\eta})$ simply singular if it is $f_{1}$-singular with $f_{1}$ defined in (1.1) and denote the set of all singular pairs $\widehat{\bf Sing}_{n,m}$.
@@
$\left(\begin{array}[]{c}\theta_{1,1}\cr\vdots\cr\theta_{n,1}\end{array}\right),\dots,\left(\begin{array}[]{c}\theta_{1,m}\cr\vdots\cr\theta_{n,m}\end{array}\right),\left(\begin{array}[]{c}1\cr\vdots\cr 0\end{array}\right),\dots,\left(\begin{array}[]{c}0\cr\vdots\cr 1\end{array}\right)$ (2.1)
@@ 2.1.3 Diophantine exponents
has solutions ${\bf q}\in\mathbb{Z}^{m}\setminus\{{\bf 0}\}$ for any $t\in\mathbb{R}_{+}$ large enough. We define the corresponding ordinary and uniform Diophantine exponents of a matrix $\Theta$ by
@@ 2.1.4 HAW property
Lastly, we briefly discuss the Hyperplane absolute winning game and the HAW property.
Hyperplane absolute game is a version of the classical two-player Schmidt’s game (see [47] for the Schmidt’s $(\alpha,\beta)$-game setup) introduced by Broderick, Fishman, Kleinbock, Reich and Weiss in [9], which provides the notion of winning stronger than the original Schmidt’s game does. We call the set Hyperplane absolute winning (HAW) if it is winning in this game. We will not need the definitions and rules of the game and thus will not define it here, just mentioning several simple properties of HAW sets which are of interest for this paper:
@@
  • A countable intersection of HAW sets is HAW.
@@ 2.2 Technical conditions on functions $f$ and $g$
Property (2.2) is natural: it is equivalent to the existence of such $K>1$ that
@@
For instance, it is satisfied by all the functions we use in Section 4. For $f$ and $g$ related via (1.2), this is equivalent to saying that
$g(CT)\geq C^{-\frac{1}{\gamma}}g(T)\,\,\,\,\text{for any}\,\,T>1\,\,\text{and for any}\,\,C\,\,\text{large enough}.$ (2.3)
The remaining stricter conditions are only used in Theorem 7 and Theorem 4 (c).
@@
$\text{There exists}\,\,\alpha>0,\,\,\text{such that}\,\,\lambda(T^{\delta})\geq\delta^{-\alpha}\lambda(T)\,\,\text{for any}\,\,\delta>1,\,\,T>0.$ (2.4)
@@
$\frac{f(T_{2})}{f(T_{1})}<\left(\frac{\log T_{1}}{\log T_{2}}\right)^{\alpha}\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,T_{2}>T_{1},$ (2.5)
where $\alpha$ is as in (2.4) (which says that $f$ locally decreases not too slowly), and
$f(T)\geq T^{-\beta}\,\,\,\,\,\,\,\,\ \text{for some}\,\,\beta>0$ (2.6)
@@
Finally, below are similar conditions needed in Theorem 4 (c).
$\frac{g(T_{2})}{g(T_{1})}<\left(\frac{\log T_{1}}{\log T_{2}}\right)^{\frac{n}{m}\alpha}\,\,\,\,\,\,\,\,\,\,\text{for any}\,\,T_{2}>T_{1},$ (2.7)
where $\alpha$ is as in (2.4) and
$g(T)\leq T^{-\frac{1}{\beta}}\,\,\,\,\,\,\,\,\ \text{for some}\,\,\beta>0.$ (2.8)